OpenStudy (anonymous):

find wronskian without solving cos(t)y''+sin(t)y'-ty=0

OpenStudy (anonymous):

I was thinking \[\cos(t)r ^{2}+\sin(t)r -t\]but I think you need to get that equation only in terms of r

OpenStudy (anonymous):

I'm now thinking you need to solve it as a Y(h) free response problem

OpenStudy (anonymous):

yeah try that cost m^2 + sint m-t=0 m1=[-sin t +sqrt(sin^2 t -4t cos t)]/2cos t m2=[-sin t -sqrt(sin^2 t -4t cos t)]/2cos t or m1=-(1/2)tam t +sqrt(sin^2 t -4t cos t) m2=-(1/2)tan t -sqrt(sin^2 t -4t cos t) W(e^-(1/2)tam t +sqrt(sin^2 t -4t cos t),e^-(1/2)tam t -sqrt(sin^2 t -4t cos t) W=[e^m1 e^ m2] [e^m1' e^m2'] you can do derivative of it

OpenStudy (anonymous):

do the derivative of y1=e^-(1/2)tam t +sqrt(sin^2 t -4t cos t) and y2=e^-(1/2)tam t -sqrt(sin^2 t -4t cos t)

OpenStudy (anonymous):

the wroskian W=[y1 y2] [y1' y2']